OpenStudy (anonymous):

How do you convert area to volume? The question I have is "Find the volume of a cube that has a total area of 96 sq. in."

OpenStudy (mathmale):

This is not really a matter of converting area to volume. Instead, you need to determine the length of a side of one of the SQUAREs that make up the total volume 96 square inches. Once you have that side length, you find the VOLUME of the cube. Hint #1: What is the formula for the area of a cube of side length x? Hint #2: What is the formula for the volume of a cube of side length x? From Hint #1, find the length of any side of the given cube.

OpenStudy (anonymous):

Is the length to one side 24 sq. in. ?

OpenStudy (anonymous):

Area of one face of the cube (6 faces) : 96/6 = 16 sq" One side of the square: V16 = 4 "

OpenStudy (mathmale):

Area of one face of the cube (6 faces) : (96 square inches)/6 = 16 square inches per side. One side of the square: Sqrt(16)=4. (Thanks to thu1935) Now you have the length of a side. What is the volume of the cube?

OpenStudy (anonymous):

Volume: (4)^3 = 4 x 4 x 4 = 64 cu"

OpenStudy (mathmale):

@CastoD33: Please make sure you take notes on the formulas for the surface area and volume of a cube; you will need these formulas later. Any questions about this particular problem?